Solve the system of equations by graphing.
y= -3x +3
6x+y=6

Answer by stanbon(75887)

Solve the system of equations by graphing.
y= -3x +3
6x+y=6
---
y = -3x + 3
y = -6x + 6
----
Substitute:
-3x + 3 = -6x + 6
3x = 3
x = 1
------
Solve for "y":
y = -3x+3
y = -3*1+3 = 0
---
Solution: (1,0)
=================
Cheers,
Stan H.
=========